OpenStudy (anonymous):

need help solving 2sqrt(3) over sqrt 10

OpenStudy (anonymous):

Is this the question? \(\Large{\frac{2\sqrt{3}}{\sqrt{10}}}\)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

I need to find the value in simpliest form

OpenStudy (anonymous):

If you get rid of the sqrt in the denominator, you have to multiply oth the top and bottom by sqrt10. Therefore, you have 2sqrt30/10. Simplifying this gives you sqrt30/5.